OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 6, find y when x = 10

OpenStudy (tkhunny):

"x varies inversely with y" MEANS \(x = \dfrac{k}{y}\). Any time you encounter such a phrase, just write it down. "q varies inversely with p" MEANS \(q = \dfrac{k}{p}\) "f varies inversely with r" MEANS \(f = \dfrac{k}{r}\) "Pittsburgh varies inversely with Cleveland" MEANS \(Pittsburgh = \dfrac{k}{Cleveland}\)